# Multimodal Fusion and Coherence Modeling for Video Topic Segmentation

### v1

@@ Appendix B Details of Annotations for the CLVTS Dataset
For the CLVTS dataset, we ensure high accuracy and reliability of annotations from three aspects. Firstly, before the actual annotation process, the annotators take two rounds of training. Each annotator needs to annotate 5 videos in each round; at the end of each round, we assess the annotation quality, provide feedback, and ensure that the annotators address the issues and understand the annotation guideline clearly at the end of training. Secondly, during annotation, to help the annotators thoroughly understand the course content, we ask the annotators to annotate topics hierarchically, that is, they label both coarse-grained topic (large topic) and fine-grained topic (small topic) boundaries while the large topic boundaries are a subset of the small topic boundaries. We take the small topic boundaries as the final topic boundary labels. Thirdly, we employ a multi-annotator strategy. All data is annotated in batches, with two annotators annotating each sample independently. The third annotator reviews the annotations by the first two annotators, rectifies errors, and provides the final annotations. After the annotations of each batch, we randomly select 5% of a batch for quality assessment. If the unacceptable rate (the proportion of the wrongly annotated topic boundaries) is less than 10%, the data are deemed satisfactory and accepted; otherwise, after communicating quality assessment results and possible reasons for errors, the annotators are requested to re-annotate the batch based on the feedback. This quality control process is repeated until the unacceptable rate is lower than 10% for all batches; in this work, we could finish quality control within 3 iterations.
@@
Table 5: Number of trainable parameters of our MMVTS models. M is the abbreviation for million.
Appendix C Training Details
@@
Our experiments are implemented with transformers package [^7] https://github.com/huggingface/transformers. For video with the number of clips greater than the max sequence length, we use sliding window and take the last clip of the prior sample as the start clip of the next sample. All supervised models use a threshold strategy, where clips with scores above a threshold 0.5 are predicted as topic boundaries. Following Gupta et al. (2023), for each video clip, we extract three visual feature types: OCR, 2D and 3D. Specifically, the OCR feature is derived by encoding the textual output obtained from the OCR API [^8] https://help.aliyun.com/zh/viapi/developer-reference/api-sy75xq of the clip’s central image. This encoding task is performed using the BERT-based sentence transformer model [^9] https://www.sbert.net/docs/sentence_transformer/pretrained_models.html, where all-mpnet-base-v2 and paraphrase-multilingual-mpnet-base-v2 models are employed for English and Chinese language experiments, respectively. The 3D feature are extracted using the same video feature extraction pipeline as in  Gupta et al. (2023), while the 2D feature is extracted by visual encoder of CLIP Radford et al. (2021) which is pre-trained to predict if an image and a text snippet are paired together. Specifically, the images from AVLecture and CLVTS are processed to extract 2D features using CLIPViT-B/16101010https://github.com/openai/CLIP and CN-CLIPViT-B/16111111https://github.com/OFA-Sys/Chinese-CLIP, respectively.
@@
After extracting the visual features, we concatenate them as shown in Eq. 16 to get $v_{i}$, which then will be fed into following projection layer. During pre-training and fine-tuning, the parameters of visual encoders are kept fixed. The learning rate is $55e-$ 5 and dropout probability is 0.1. AdamW Loshchilov and Hutter (2017) is used for optimization. The batch size is 8 and the epoch for pre-training and fine-tuning is 1 and 5, respectively. The loss weight $\alpha$ and $\gamma$ for $l_{cma}$ is 0.5, $\beta$ and $\sigma$ for $l_{balance}$ is 1.0 when MoE is in multimodal fusion layer. $\theta$ for $l_{mcssl}$ is 0.5. $k_{1}$ and $k_{2}$ of Eq. 14 is 1 and 3 following Yu et al. (2023). We select one multimodal fusion layer to comprehensively compare different types of fusion structure. We also investigate the performance of different number of multimodal fusion layer in Figure 3. As for the MoE layer in multimodal fusion layer, we choose 4 candidate experts and activate 2 experts for each input feature, the intermediate size of expert is 3072. Therefore, the total number of trainable parameters is shown in Table 5.
@@
Appendix D Evaluation Metrics
F1 is a metric used to evaluate the accuracy of text topic segmentation Lukasik et al. (2020); Zhang et al. (2021), which focus on the performance of exact matching of the positive class and balances precision and recall rate.
BS@k Gupta et al. (2023) is the average number of predicted boundaries matching with the ground truth boundaries within a k second interval, which can be regard as the recall rate based on fuzzy matching.
F1@k denotes the F1 score calculated based on matching predicted and ground truth boundaries within k seconds. Considering subjectivity and uncertainty in VTS annotations, we introduce F1@k as a supplement to BS@k to enabling a more comprehensively assessment of model performance in dealing with ambiguously defined boundaries.
mIoU is commonly used in video action segmentation Zhou et al. (2018) and video scene segmentation Mun et al. (2022). While the F1, BS@k and F1@k focus on the accuracy of positive predictions (either exact match or fuzzy match), the mIoU measures the overlapping area between predicted and ground truth segments, providing a generalized assessment of how well the model’s predicted segments match the actual segments in the segmentation task.
Our implementation of BS@30 draws upon the code publiced by Gupta et al. (2023) [^12] https://github.com/Darshansingh11/AVLectures/, while the approach to implement mIoU is guided by Mun et al. (2022) [^13] https://github.com/kakaobrain/bassl. We have relied on the scikit-learn package [^14] https://scikit-learn.org/stable/ to compute F1, following the implementation by Yu et al. (2023). The definitions provided in the aforementioned sources also inform our implementation of F1@k.
Figure 3: Avg performance of fine-tuning model w/ or w/o pre-training and coherence modeling tasks on AVLecture with different numbers of multimodal fusion layers, where the fusion type is Co-Attention with MoE.
Appendix E Artifact Use Consistent With Intended Use
@@
Appendix F Fine-tune Llama-3-8B for Video Topic Segmentation
Considering the computational complexity and data volume, we employ LoRA Hu et al. (2021) for fine-tuning Llama-3-8B [^15] https://huggingface.co/meta-llama/Meta-Llama-3-8B with a maximum sequence length of 2048. Our training configuration includes a batch size of 32 and a total of 6 epochs, utilizing a learning rate of $55e-$ 5 and cosine scheduler Loshchilov and Hutter (2016). Table 7 shows two prompts we used to fine-tuning Llama-3-8B on text data of VTS. We first try the Generative type that Yu et al. (2023) has experimented with due to the better zero-shot and one-shot performance than the Discriminative type. However, compared with Longformer, the Avg score of Generative prompting Llama-3-8B is lower 8.3 and 6.27 on AVLecture and CLVTS respectively. We suspect that this is due to the issue of sparse labels in binary classification, leading us to refine the Discriminative into the Discrete prompt as outlined in the Table 7. Although the $F_{1}$ on AVLecture and CLVTS is much lower (the prediction boundary of the model usually has a clip offset), the Avg score of Llama-3-8B on AVLecture is merely 2.26 points behind LongFormer. Moreover, Llama-3-8BDiscrete secures the highest Avg score on CLVTS, surpassing LongFormer by 4.53 points. which suggests that the Discrete prompt is more suitable for stimulating Llama-3-8B and future work can further explore how to better use large language models for VTS.
Type Prompts for text topic segmentation
@@ Appendix F Fine-tune Llama-3-8B for Video Topic Segmentation
Table 7: Prompts for fine-tuning Llama-3-8B on text data of video topic segmentation. $n$ denotes the number of sentences and $s_{i}$ denotes $i$-th sentence in the document.
Appendix G Performance of MFLs across varied layer numbers
@@
Utilizing Co-Attention with Mixture of Experts (MoE) as the multimodal fusion layer, we investigated the Avg performance of models featuring various numbers of Multimodal Fusion Layers (MFLs) on AVLecture, as depicted in Figure 3. The findings indicate that when the model is fine-tuned directly on labeled data, a single MFL yields the best performance, surpassing the no-layer configuration by 2.34 points. However, adding more layers leads to diminished performance and training instability, particularly noticeable with three layers. By incorporating a pre-training phase followed by fine-tuning through a coherence modeling task, performance enhancements of 4.0, 2.87, 2.37, and 13.51 points are observed, respectively. These results clearly demonstrate the substantial benefits of our pre-training and coherence modeling strategies in boosting the model’s effectiveness. Notably, they facilitate the convergence of a model with three layers, achieving optimal results that marginally exceed the performance of a single-layer model by 0.1 points.

### v2

@@ A.1 Data Collection and Annotation
Thirdly, we employ a multi-annotator strategy. All data is annotated in batches, with two annotators annotating each sample independently. The third annotator reviews the annotations by the first two annotators, rectifies errors, and provides the final annotations. After the annotations of each batch, we randomly select 5% of a batch for quality assessment. If the unacceptable rate (the proportion of the wrongly annotated topic boundaries) is lower than 10%, the data are deemed satisfactory and accepted; otherwise, after communicating quality assessment results and possible reasons for errors, the annotators are requested to re-annotate the batch based on the feedback. This quality control process is repeated until the unacceptable rate is lower than 10% for all batches. In this work, we finished quality control of all data within 3 iterations.
A.2 Dataset Analysis
To evaluate the inter-annotator agreement on VTS annotations on the CLVTS dataset, following Shou et al. (2021), we compute the $F_{1}@k$ score (defined in Appendix C) based on the absolute distance between two topic boundary sequences, varying the threshold $k$ from 0 to 8 seconds with a step size of 2 seconds, where 8 seconds are approximately the average duration of a video clip. By averaging the F1 scores across all three pairs of annotated topic boundaries from three annotators on the same video, we obtain the consistency score. The more similar the annotations from all annotators on the same video are, the higher the consistency score is. Figure 2 shows that the consistency scores of the majority of videos in our CLVTS dataset exceed 0.5, indicating a decent degree of consensus for VTS annotations Shou et al. (2021).
Table 1 compares our CLVTS dataset against existing VTS datasets. Both AVLecture and CLVTS are sourced from educational contexts, where VTS significantly enhances learning experiences. Table 1 highlights that CLVTS features a higher average number of topics per video. Among annotated videos in CLVTS, 47% are presentations showing slides, 34% are blackboard demonstrations, and 19% are miscellaneous types. We also collect 1027 hours of unlabeled videos from the same sources for pre-training. Figure 2 and 2 show a diverse distribution of video durations and topic durations and a broad spectrum of subjects in labeled CLVTS.
A.3 Ethical Considerations
The dataset used in this research is strictly for academic and non-commercial purposes. We implemented several measures to ensure compliance with ethical standards, as follows.
  • Data Transparency and Anonymization. We only provide ASR transcripts after rigorous text anonymization processes, visual features of video clips, our annotations, and links to the original videos, to ensures transparency regarding the data sources and their usage while maintaining anonymity.
  • Data Access Compliance. To further ensure ethical use of the dataset, we require researchers to contact us via emails to confirm their compliance with ethical guidelines and the conditions outlined in our data usage declaration, before granting them access to the dataset. This procedure includes ensuring that they are aware of and adhere to the Personal Information Protection Law (PIPL) and any relevant legal frameworks regarding personal data usage.
  • Authorization. Any personal data should be used only with express authorization, ensuring lawful and fair processing in accordance with applicable laws.
@@ A.3 Ethical Considerations
Table 5: The number of trainable parameters of the baseline LongFormercssl and variants of our MMVTS models. The model names conform to the model names in Table 2.
Appendix B Implementation Details
We partition the labeled data within AVLecture and CLVTS into 70% for training, 10% for validation, and 20% for testing, respectively. The unlabeled data of AVLectures and CLVTS are used for pre-training for each dataset, respectively.
Our experiments are implemented with the transformers package [^7] https://github.com/huggingface/transformers. We use the same maximum sequence length 2048 as in Yu et al. (2023) for a fair comparison with the text-only models. All results are the mean values over three runs with different random seeds.
For video with the number of clips greater than the max sequence length, we use sliding window and take the last clip of the prior sample as the first clip of the next sample. All supervised models use a threshold strategy, where clips with scores above a threshold 0.5 are predicted as topic boundaries.
@@ Appendix B Implementation Details
Following Gupta et al. (2023), for each video clip, we extract three visual feature types: OCR, 2D and 3D. Specifically, the OCR features are derived by encoding the textual output obtained from the OCR API [^8] https://help.aliyun.com/zh/viapi/developer-reference/api-sy75xq of the clip’s central image. Encoding the textual output from OCR is performed using the BERT-based sentence transformer model [^9] https://www.sbert.net/docs/sentence_transformer/pretrained_models.html, where all-mpnet-base-v2 and paraphrase-multilingual-mpnet-base-v2 models are employed for experiments on the English and Chinese datasets, respectively. The 3D features are extracted using the same video feature extraction pipeline as in  Gupta et al. (2023). The 2D features are extracted by sampling three frames from each clip, subsequently encoding these frames with visual encoder, and applying max pooling. Specifically, we choose the visual encoder of CLIP Radford et al. (2021), which is pre-trained to predict if an image and a text snippet are paired together. The images from AVLecture and CLVTS are processed to extract 2D features using CLIPViT-B/16101010https://github.com/openai/CLIP and CN-CLIPViT-B/16111111https://github.com/OFA-Sys/Chinese-CLIP, respectively.
@@
After extracting the visual features, we concatenate them as shown in Eq. 16 to get $v_{i}$, which will then be fed into the following projection layer. During pre-training and fine-tuning, the parameters of visual encoders are kept frozen. The learning rate is $55e-$ 5 and dropout probability is 0.1. AdamW Loshchilov and Hutter (2017) is used for optimization. The batch size is 8 and the epoch for pre-training and fine-tuning is 1 and 5, respectively. The loss weight $\alpha$ and $\gamma$ for $l_{cma}$ is 0.5, $\beta$ and $\sigma$ for $l_{balance}$ is 1.0 when MoE is in Multimodal Fusion Layers, $\theta$ for $l_{mcssl}$ is 0.5. $k_{1}$ and $k_{2}$ of Eq. 14 are 1 and 3, following Yu et al. (2023). We comprehensively compare different types of fusion structure using one Multimodal Fusion Layer, then use the best performing fusion structure for the remaining experiments. We also investigate the impact of different numbers of Multimodal Fusion Layers in Figure 3. As to the MoE layer in Multimodal Fusion Layers, we choose 4 candidate experts and activate 2 experts for each input feature. The intermediate size of expert is 3072. The total number of trainable parameters is shown in Table 5.
@@
Appendix C Evaluation Metrics
F1 is a metric used to evaluate the accuracy of text topic segmentation Lukasik et al. (2020); Zhang et al. (2021). It focuses on the performance of exact matching of the positive class and balances the precision and recall rates.
BS@k Gupta et al. (2023) is the average number of predicted boundaries matching with the ground truth boundaries within a k-second interval, which can be considered as the recall rate based on fuzzy matching.
F1@k denotes the F1 score calculated based on matching predicted boundaries and ground truth boundaries within k seconds. Considering subjectivity and uncertainty in VTS annotations, we introduce F1@k as a supplement to BS@k to enabling a more comprehensive assessment of model performance in dealing with ambiguous (uncertain) boundaries.
mIoU is commonly used in video action segmentation Zhou et al. (2018) and video scene segmentation Mun et al. (2022). While the F1, BS@k and F1@k metrics focus on the accuracy of positive predictions (either exact match or fuzzy match), the mIoU metric measures the overlapping area between predicted segments and ground truth segments, hence providing a generalized assessment of how well the model’s predicted segments match the actual segments on the segmentation task.
Our implementation of BS@30 draws upon the code published by Gupta et al. (2023) [^12] https://github.com/Darshansingh11/AVLectures/, while the approach to implement mIoU is guided by Mun et al. (2022) [^13] https://github.com/kakaobrain/bassl. We have relied on the scikit-learn package [^14] https://scikit-learn.org/stable/ to compute F1, following the implementation by Yu et al. (2023), to ensure fair comparisons. The definitions provided in the aforementioned sources also inform our implementation of F1@k.
Figure 3: Performance (Avg) of fine-tuning our MMVTS model on AVLecture test set, w/o pre-training and multimodal coherence modeling tasks (denoted by w/o pre and coh) and w/ pre-training and multimodal coherence modeling tasks (denoted by w/ pre and coh), with different numbers of Multimodal Fusion Layers, where the fusion structure is Co-Attention with MoE.
Appendix D Artifact Use Consistent With Intended Use
@@
(a) Introduction to Nuclear and Particle Physics
(b) Latency and Throughput
(c) Quiz Review From Optional Problem Set 8
(d) Quantum Mechanics
(e) Attention
(f) Vision
Figure 4: Video Topic Segmentation examples for six lecture videos from AVLecture. (a)-(b) are of the Slides mode, (c)-(d) are Blackboard mode and (e)-(f) are Mixed mode. GT denotes ground truth, V refers to the visual-only baseline BaSSL which only uses visual features, T denotes the text-only LongFormer which only relies on the text modality, and V+T signifies our MMVTS model that integrates information from both visual and text modalities. Unsup denotes the unsupervised baseline UnsupAVLS, which also integrates visual and text features.
Appendix E Fine-tune Llama-3-8B for Video Topic Segmentation
Considering the computational complexity and the data volume, we employ LoRA Hu et al. (2021) for fine-tuning Llama-3-8B [^15] https://huggingface.co/meta-llama/Meta-Llama-3-8B on the training set of each AVLecture and CLVTS datasets, with a maximum sequence length of 2048. With LoRA, the number of trainable parameters is 3 million. Our training configuration includes a batch size of 32 and a total of 6 epochs, utilizing a learning rate of $55e-$ 5 and cosine scheduler Loshchilov and Hutter (2016). Table 7 shows two prompts we used to fine-tuning Llama-3-8B on the text modality of VTS data. We first evaluate the Generative type prompt that Yu et al. (2023) has experimented with, due to its better zero-shot and one-shot text topic segmentation performance than the Discriminative type. However, compared with Longformer which also uses the textual modality, the Avg score of fine-tuning Llama-3-8B with the Generative prompting strategy (denoted by Llama-3-8BGenerative) is 8.3 and 6.27 absolutely worse than those from Longformer on AVLecture and CLVTS, respectively. We suspect that this is due to the inherent issue of sparse labels in binary classification for VTS, posing challenges to applying the large language model (LLM) Llama-3-8B in a generative manner. Inspired by this hypothesis, we refine the Discriminative prompt into the Discrete prompt for Llama-3-8B, as shown in Table 7. The results in Table 2 show that on Avg, Llama-3-8BDiscrete consistently and markedly outperforms Llama-3-8BGenerative on both AVLecture and CLVTS, by 6.04 and 10.8 absolute points, respectively. These results suggests that the Discrete prompt is more suitable for prompting Llama-3-8B for the VTS task than the Generative prompt.
However, with either Generative or Discrete prompt, Llama-3-8BDiscrete does not provide consistent advantages over the text-only small model Longformer. Table 2 shows that $F_{1}$ results of Llama-3-8BDiscrete on AVLecture and CLVTS are much worse than those from Longformer (we find that the prediction boundary of Llama-3-8B usually has a clip offset). The Avg score of Llama-3-8BDiscrete on AVLecture is 2.26 absolute worse than LongFormer, although Llama-3-8BDiscrete achieves the best Avg score on CLVTS, surpassing LongFormer by 4.53 absolute. Future work could further explore how to better use LLMs for the VTS task.
Type Prompts for text topic segmentation.
@@ Appendix E Fine-tune Llama-3-8B for Video Topic Segmentation
Table 7: Our designed prompts for fine-tuning Llama-3-8B on the text data for video topic segmentation. $n$ denotes the number of sentences and $s_{i}$ denotes the $i$-th sentence in the document.
Appendix F More Analysis of the Baseline Results
Considering the baseline results in Table 2, the only seemingly similar scores between the unsupervised method and supervised baselines are mIoU scores, which are attributed to the leakage of the ground-truth topic number. If we use the topK probability to determine predictions, where K for each sample is the ground-truth topic number, mIoU and BS@30 of $SWST_{seq}$ (Row 7) will be 4.4 and 19.05 points higher than UnsupAVLS, respectively. However, this evaluation is not reasonable, as models should not disclose the ground-truth topic number during testing. Therefore, we opted for the threshold-based evaluation strategy commonly used in classification tasks
Appendix G Performance of Multimodal Fusion Layers with Varied Numbers of Layers
@@
Using Co-Attention with MoE as the architecture of Multimodal Fusion Layers, we investigate the Avg performance of models featuring various numbers of Multimodal Fusion Layers (MFLs) on the AVLecture data set, as depicted in Figure 3. We find that directly fine-tuning our MMVTS model, without pre-training nor the two auxiliary tasks for coherence modeling, a single MFL yields the best performance, surpassing the no-layer configuration by 2.34 points. Adding more layers leads to degraded performance and training instability, particularly noticeable with three layers.
By incorporating the pre-training phase followed by fine-tuning with coherence modeling tasks, Avg performance enhancements of 4.0, 2.87, 2.37, and 13.51 points for 0, 1, 2, 3 MFL layers are observed, respectively. These results clearly demonstrate the substantial benefits from our pre-training and coherence modeling strategies in boosting the model’s performance. Notably, our pre-training and coherence modeling reaches the convergence of the model with three MFL layers, achieving results that marginally exceed the performance of a single-layer model by 0.1 points.
Model Modality AVLecture CLVTS
Longformer T 62.52 46.81
Llama-3-8BDiscrete T 60.26 51.34
MMVTS Model (Ours) V+T 68.61 51.27
V+T+A1 67.49 51.98
V+T+A2 69.36 52.97
Table 8: The Avg score of integration of audio information into the MMVTS model with Co-Attn and MoE architecture, on AVLecture and CLVTS test sets. V+T+A1 notes that during fine-tuning phase, the audio features are incorporated into the cross-modality alignment loss lcma, while V+T+A2 does not involve audio features in lcma.
Appendix H Qualitative Analysis of Video Topic Segmentation Examples
We present the topic-segmented outputs for six lecture videos from three presentation modes including slides, blackboard, and mixed, in Figure 4. Relying solely on the visual modality, segmentation points are predominantly identified through the superficial cues associated with visual transitions. In contrast, using only the textual modality, topic boundaries are discerned based on semantic content; however, this approach has limitations, such as missing topic boundaries or the accumulation of topics. The integration of the visual modality with the text modality offers complementarity of information, thereby improving the overall VTS performance. The case studies in Figure 4 help illustrate the importance of multimodal fusion for the VTS task regardless of the video presentation modes.
Appendix I Integration of Audio Information
We use a pre-trained Whisper-small model [^16] https://github.com/openai/whisper/ with 244 million parameters and use the Whisper encoder to extract audio features for video clips (the parameters of the Whisper audio encoder are frozen). Similar to visual features, we perform maximum pooling on the encoded audio features for each clip to obtain the clip-level audio features. Then, we feed the audio features into a learnable projection layer and then through the Multimodal Fusion Layers, and then concatenate them with visual features and text features for the final topic boundary prediction.
We initialize the parameters from pre-trained visual and textual MMVTS model with Co-Attn and MoE. During fine-tuning, we compare two configurations. The first configuration is denoted by V+T+A1, which incorporates the audio feature into the cross-modality alignment loss lcma. The second configuration is denoted by V+T+A2, which does not involve audio features in lcma. The pairwise cross-modality alignment loss weight is set to 0.33 in V+T+A1.
The results are shown in Table 8. As can be seen from the table, integration of audio features by V+T+A1 improves the Avg score by 0.71 on the CLVTS test set while decreasing the Avg socre by 1.12 on the AVLecture test set. However, we find that excluding audio from the lcma (V+T+A2) results in an absolute improvement of 0.75 and 1.7 in the Avg score on AVLecture and CLVTS test sets, respectively, which suggests that audio features can contribute to more accurate topic boundary prediction under certain conditions, but their role in modality alignment needs to be treated with caution and demands further exploration. We sample some videos and observe that the cosine similarity between audio features of adjacent clips shows relatively small differences, ranging from 0.97 to 0.98, while visual features between adjacent clips exhibit larger differences, ranging from 0.88 to 0.97. This disparity might complicate the task of simultaneously aligning text, visual, and pooled clip-level audio features. Future research could explore more nuanced integration of audio features to provide supplementary paralinguistic information, such as pitch, energy, and pause duration. Alternatively, direct use of audio and visual information for VTS may bypass the ASR step altogether (as ASR is used to obtain the text modality).